OpenStudy (anonymous):

trendy t-shirts has decided to manufacture a new design. it will cost $400 plus $3 per shirt to produce them, and trendy plans to spend $5000on advertising. the shirts will sell for $12 each A. how many shirts must be sold to break even (total cost equals income from sales)? B . what is trendy's income (or total cost) for that many shirts?

OpenStudy (anonymous):

let x be the no. of t-shirts sold Then cost price=400+3x+5000 $ selling price=12x $ now 12 x=400+3x+5000 calculate x to find the income( or total cost) put in 12x